### v2

# Towards Reliable Latent Knowledge Estimation in LLMs: Zero-Prompt Many-Shot Based Factual Knowledge Extraction

Abstract.
In this paper, we focus on the challenging task of reliably estimating factual knowledge that is embedded inside large language models (LLMs). To avoid reliability concerns with prior approaches, we propose to eliminate prompt engineering when probing LLMs for factual knowledge. Our approach, called Zero-Prompt Latent Knowledge Estimator (ZP-LKE), leverages the in-context learning ability of LLMs to communicate both the factual knowledge question as well as the expected answer format. Our knowledge estimator is both conceptually simpler (i.e., doesn’t depend on meta-linguistic judgments of LLMs) and easier to apply (i.e., is not LLM-specific), and we demonstrate that it can surface more of the latent knowledge embedded in LLMs. We also investigate how different design choices affect the performance of ZP-LKE. Using the proposed estimator, we perform a large-scale evaluation of the factual knowledge of a variety of open-source LLMs, like OPT, Pythia, Llama(2), Mistral, Gemma, etc. over a large set of relations and facts from the Wikidata knowledge base. We observe differences in the factual knowledge between different model families and models of different sizes, that some relations are consistently better known than others but that models differ in the precise facts they know, and differences in the knowledge of base models and their finetuned counterparts. 111Code available at: https://github.com/QinyuanWu0710/ZeroPrompt_LKE
1. Introduction
Conversational chatbots (e.g., OpenAI’s ChatGPT) built around large language models (e.g., OpenAI’s GPT) are increasingly being used for a variety of information retrieval tasks such as searching for information or seeking recommendations related to real-world entities like people or places (Wu et al., 2023; Zhu et al., 2023). A worrisome concern in such scenarios is the factual correctness of information generated by the LLMs (Peng et al., 2023; Hu et al., 2023b; Snyder et al., 2023; Yao et al., 2023; Ji et al., 2023; Zhang et al., 2023; Wang et al., 2023; Kalo and Fichtel, [n. d.]; Cao et al., 2021; Youssef et al., 2023).
The latent knowledge estimation problem: To avoid making false assertions about a real-world entity, an LLM first needs to have factual (true) knowledge about the entity. Given a prompt like “Einstein was born in the year”, LLMs may generate both the correct answer (“1879”) and wrong answers (e.g., “1878” or “1880”) with some probabilities. If an LLM knows the fact, one can hope that the probability with which it would generate the correct answer would be much higher than the wrong answers (Jiang et al., 2021). As LLMs are typically pretrained over a Web corpus (including Wikipedia data) with millions of facts about real-world entities, they have the opportunity to learn factual knowledge about our world and latently embed the knowledge in their parameters. But, how can we estimate the extent of LLMs’ knowledge of real-world facts?
Reliability of latent knowledge estimates: Following (Petroni et al., 2019), many prior works (Jiang et al., 2020; Bouraoui et al., 2020; Shin et al., 2020) represent factual knowledge in the form of triplets , where the subject has a relation of type with the object (e.g., ). The central challenge of latent knowledge estimation is to infer given and by only using information extracted from the LLM. Typically, the inference relies on probing the LLM with prompt templates, , constructed to communicate the information of and and analyzing the generated responses for presence of (see Figure 1). Current approaches (Hu et al., 2023a; Kalo and Fichtel, [n. d.]; Jiang et al., 2021, 2020; Shin et al., 2020; Sun et al., 2023) allow unrestricted choice of prompt templates with few well-defined rules (see Figure 2). As a result, they are vulnerable to prompt engineering and prompt hacking, which raises serious concerns about the reliability of their estimates (Cao et al., 2021). Against this background, in this paper, we make four primary contributions:
1. A structured reliable latent knowledge estimator (LKE) based on zero prompting (ZP): Our latent knowledge estimator, called ZP-LKE, is based on the following simple, yet novel and powerful insight. Rather than engineer input prompts that best communicate to an LLM, we let the LLM infer by simply providing multiple examples of pairs that share the same relation (see Figure 2). The key distinguishing feature of ZP-LKE is its adherence to zero prompting, i.e., the input is highly structured and contains no prompt tokens outside other similarly related pairs. Thus, ZP-LKE avoids the reliability risks associated with prompt engineering such as side-channels and over-fitting. (We discuss these reliability concerns in Section 2.1.)
| Prompt based LKEs (ZS-LKE & FS-LKE) | Zero-Prompt based LKE (ZP-LKE) | ||
| Input | Output: Next tokens | Input | Output: Next tokens |
| Peter Grünberg was born in | Munich in 1939 and obtained | Albert Einstein 1879, Peter Grünberg | 2007 Mohammed Hanif |
| Albert Einstein was born in 1879, Brian Kobilka was born in 1955, Peter Grünberg was born in | 1939 and has a birthday | Albert Einstein 1879, Brian Kobilka 1955, Stefan W. Hell 1962, Ivan Pavlov 1849, Peter Grünberg | 1939, Albert Szent- |
| Albert Einstein was born in 1872, Brian Kobilka was born in 1965, Peter Grünberg was born in | 1939 and has a birthday | Albert Einstein 1872, Brian Kobilka 1927, Stefan W. Hell 1962, Ivan Pavlov 1878, Peter Grünberg | 2007, Albert Szent- |
| Author-1 was born in 1872, Author-2 was born in 1965, Peter Grünberg was born in | 1939 and has a life expect | Author-2 1879, Author-3 1955, Author-4 1962, Author-1 1849, Peter Grünberg | 2022, Rossitza |
2. ZP-LKE requires many-shots and is fundamentally different from few-shot prompting: A recent work (Kalo and Fichtel, [n. d.]) shows that few-shot prompting (FS-LKE in Figure 2) can yield improved knowledge estimates compared to zero-shot prompting (ZS-LKE in Figure 2) by providing the LLM with examples to format generated answers. In contrast, ZP-LKE also uses examples to effectively communicate the question at hand. The different modes in which examples are being used in FS-LKE and ZP-LKE is illustrated further in Table 1. Adding a few examples to the zero-shot prompt ”Peter Grünberg was born in” results in the model generating the correct answer ”1939” right-away. However, it does not appear to matter whether the examples provided are reflecting correct information or information about subjects known to the LLM (consistent with our hypothesis that the examples are being used to infer answer format). In contrast, Table 1 suggests that for ZP-LKE not only is the number of examples needed larger, but it also matters whether they are correct and known to the LLM (consistent with our hypothesis that examples are being used to infer the question). The two distinct ways in which examples are being used by FS-LKE and ZP-LKE map well to the dual modes of in-context learning namely, task recognition and task learning, respectively, that have been identified in a recent work(Pan, 2023).
We systematically investigate how factors such as how many examples are provided in an ZP-LKE, whether some of those examples are unknown to the model or simply incorrect, as well as how the examples are ordered affect knowledge estimation. We find that ZP-LKE requires many-shots, which make it relatively robust to unknown examples, but ZP-LKE remains vulnerable to incorrect examples. Our findings represent a nuanced exploration of in-context learning(Brown et al., 2020), where the dominant learning mode is task learning rather than task recognition.
3. ZP-LKE significantly outperforms previous prompt-based approaches across different open-source models and different types of factual relations: We empirically compared the performance of ZP-LKE against prior approaches that relied on a variety of human-generated prompts (HGPs) as well as machine-mined prompts (MMPs)(Jiang et al., 2020). Across a large set of facts spanning different types of relations from the widely-used T-Rex dataset (Elsahar et al., 2018), we find that ZP-LKE improves the fraction of facts accurately extracted from four open-source models by an average of 35% for HGPs (from 0.45 to 0.61) and 90% for MMPs (from 0.32 to 0.61). These performance gains of ZP-LKE arise from a better comprehension of the question as well as the expected answer format. To quantify the performance gains from only better question comprehension, we propose in Section 2.2 a multiple-choice testing that accounts for answer formats. We find that ZP-LKE still outperforms existing approaches by an average of 9.41% for HGPs (from 0.71 to 0.78) and 57% for MMPs (from 0.50 to 0.78), with improvements for specific relations like ”position played on team/specialty” varying from 152% for HGPs (from 0.17 to 0.43) and 310% for MMPs (from 0.10 to 0.43). Thus, ZP-LKE represents a better way to retrieve knowledge stored internally within an LLM, surpassing the model’s ability to follow instructions in prompt templates.
4. Being model-agnostic, ZP-LKE enables a systematic comparison of latent knowledge of open source LLMs at scale: In contrast to prompt-based LKEs (Jiang et al., 2020; Shin et al., 2020), which are tailored to specific relations and models, ZP-LKE creates a single input to test for facts pertaining to a relation that can be used flexibly for any model. This simplicity and versatility allows for cross-LLM comparisions of factual knowledge. Using ZP-LKE, we evaluated the knowledge of 49 open-source LLMs from various families like Llama(2), Gemma, Mistral, OPT, and Pythia. These models vary in size and were tested with and without instruction-finetuning on 50 different relations and 20,000 facts from Wikidata. We found that models from families such as Llama2, Mistral, and Gemma, as well as larger models, know more facts. Models within the same family differ in the specific facts they know, even if trained on the same data. Additionally, instruction fine-tuning reduces the amount of factual knowledge that can be extracted from these models. Our findings will likely be of interest to developers that wish to train models with lots of embedded factual knowledge.
Related Work: Researchers have proposed several approaches (Youssef et al., 2023) to estimate latent knowledge from LLMs, which can be categorized into two main methods: (i) Model-internals based approaches: These methods use various internal aspects of the LLM, such as attention maps (Wang et al., 2020), activation functions (Burns et al., 2022), or model parameters (Kazemnejad et al., 2023), to determine whether factual information can be extracted from the model. (ii) Model-responses-based approaches, which are generally applicable to a wide range of LLM models and there are two key parts of the model-responses-based approach: constructing the input and evaluating the output from the LLM.
Input construction: There are different prompting techniques to verify if a target fact is stored in the model. These prompt-based methods differ in their choice of prompts, which can be divided into human-generated prompts (HGPs) (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Petroni et al., 2019; Jiang et al., 2021; Newman et al., 2022; Jiang et al., 2020; Kalo and Fichtel, [n. d.]) and machine-mined prompts (MMPs) (Shin et al., 2020; Jiang et al., 2020). All the prompting-based methods try to find the best template for the question that the model can comprehend best; however, the optimization of the prompting template can be unreliable (Zamfirescu-Pereira et al., 2023; Arora et al., 2023; Sclar et al., 2023; Cao et al., 2021). Instead of finding the best comprehensive template, our approach proposes using structured triplet examples to communicate and prob the tested fact, uncover deeper relations and knowledge in the LLM. This method communicates the question through in-context examples, a strategy that, to our knowledge, has not been explored before. A similar approach is to use few-shot prompting. For example, (Kalo and Fichtel, [n. d.]) first found the best prompt template and then applied few-shot prompting to guide and limit the model’s response format. However, this approach still holds the limitation of template searching and relies on the model’s comprehension of the template, which is fundamentally different from our approach.
Output evaluation: The evaluation methods of early works are LLM specific, limiting the evaluated objects to single-token outputs (Bouraoui et al., 2020; Jiang et al., 2020; Shin et al., 2020; Petroni et al., 2019). More recent works evaluate the generation by checking the next generated tokens to see whether the potentially multi-token ground truth appears in the generated tokens (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Jiang et al., 2021; Newman et al., 2022; Kalo and Fichtel, [n. d.]). However, the final performance is significantly influenced by the choice of , and the generation quality also relies heavily on various sampling parameters, which introduces uncertainty (Lin et al., 2023). In order to respond more fundamentally to the model’s level of knowledge, without focusing on metrics such as the fluency of generation, we constructed a multiple-choice dataset with 100 unique possible choices for each evaluated fact and judged whether or not the model knew the fact by comparing the relative probabilities of these 100 objects.
Factual knowledge datasets: Different from existing knowledge evaluation benchmarks like TruthfulQA (Lin et al., 2021) and MMLU (Hendrycks et al., 2020), which already provide templates of questions, our approach considers facts from existing knowledge graphs for performing knowledge estimation of LLMs. As a test bed (Elsahar et al., 2018; Hu et al., 2023a; Sun et al., 2023; Petroni et al., 2019; Zhu and Li, 2023; Kryściński et al., 2019), we utilize knowledge graphs, allowing our method to be applied to any knowledge graph database without additional effort.
2. Designing Reliable LKEs
Today, there exist many general-purpose as well as domain-specific factual knowledge bases that contain a very large number (millions to billions) of facts. The facts can be encapsulated as triplets, represented as subject (), relation (), object (). These triplets offer a general way to represent factual knowledge about real-world entities in knowledge graphs or other structured knowledge bases. The goal of latent knowledge estimation is to infer what fraction of the facts are known to an LLM. We call methods that estimate the amount of latent knowledge inside an LLM latent knowledge estimators (LKEs).
2.1. Reliability concerns with existing LKEs
Existing approaches to estimating latent knowledge in LLMs use various factual knowledge tests. We identify several reliability concerns (RCs) with current designs that motivate our new LKE design. While some related works address some of these concerns, none have comprehensively solved all the issues (Kalo and Fichtel, [n. d.]; Cao et al., 2021).
RC 1. Reliance on unrestricted prompt engineering: Many past works have attempted to use test prompts without any restrictions, including both human-generated or machine-mined prompts (Jiang et al., 2020; Zamfirescu-Pereira et al., 2023; Arora et al., 2023; Sclar et al., 2023). They typically intersperse the subject and object between additional relationship context-communicating tokens. Some analyze the performance of a variety of prompts and then pick the best-performing or use an ensemble of the best-performing prompts (Jiang et al., 2020; Newman et al., 2022; Fernando et al., 2023). However, unrestricted prompt engineering risks introducing side-channels and over-fitting. First, the generated prompts, particularly those that are machine-mined, may include tokens that can implicitly or explicitly introduce additional (side-channel) information that makes it easier to answer the question. As a specific example, in a prior work (Jiang et al., 2020), for the relation “position held”, the prompt “ has the position of ” performed worse than “ is elected ”. But, note that the second prompt potentially introduces a side-channel: it implicitly rules out answer choices for unelected positions like Professor and favors elected positions like President. Second, selecting from an unbounded number of potential prompt choices raises concerns about the complexity of LKEs (the size of the set of all considered prompts) and the associated risks of over-fitting, which in turn affect the reliability of estimates.
RC 2. Reliance on LLMs’ meta-linguistic judgments: Prior works used prompt templates with instructions (Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Petroni et al., 2019; Jiang et al., 2021; Newman et al., 2022; Jiang et al., 2020; Cao et al., 2021; Kalo and Fichtel, [n. d.]; Youssef et al., 2023; Zhao et al., 2024) for communicating the question as well as the expected format of answers. But, the scores (estimates) resulting from such prompt-based testing conflate an LLM’s latent knowledge of the facts with the LLM’s meta-linguistic judgments, i.e., the LLM’s ability to comprehend the prompt, understand the question embedded within the prompt and output the answer in some expected format (Hu and Levy, 2023). The impact on meta-linguistic judgments can be seen from the fact that multiple semantically-equivalent prompts result in different responses from an LLM and thereby, different estimates of latent knowledge (Hu and Levy, 2023).
RC 3. Reliance on LLM-specific prompts: Many prior works (Petroni et al., 2019; Jiang et al., 2020; Shin et al., 2020) limit the choice of facts that can be used in tests to those where the surface form of the objects () is represented by a single token by the LLM’s tokenizer. Even though some works are able to evaluate multiple-token objects, prompt-based approaches need careful prompt engineering for different LLMs to get the best prompt template (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Jiang et al., 2021; Newman et al., 2022; Kalo and Fichtel, [n. d.]), which makes it hard to estimate and compare factual knowledge across a large number of LLMs and the prompt optimisation would be very expensive and inefficient for large models.
Motivated by the above, we derive the following three design principles (DPs) for LKEs. A reliable LKE design should:
DP1: limit prompt hacking to avoid over-fitting & side-channels.
DP2: minimize reliance on meta-linguistic prompts.
DP3: avoid LLM-specific prompts.
2.2. A new Zero- Prompt based LKE (ZP-LKE)
Our goal is to estimate whether an LLM knows a fact . The challenge is to probe the LLM and evaluate its responses in a way compatible with the design principles defined in Section 2.1.
The key idea here is to eliminate prompts meant to capture the relation (zero-prompt) and instead rely on examples of similarly related pairs to probe the internal knowledge. LLMs have been shown to exhibit In-Context Learning (ICL) abilities (Brown et al., 2020) that allow them to infer and extrapolate patterns in their inputs. We leverage this ability to communicate information about relation without additional instructions to the LLMs (DP1 and DP2) by providing it with a list of facts based on .
Example 0.
Assume that we want to probe for whether an LLM knows the fact Einstein, birth-year, 1879 . We can use other facts for the birth-year relation such as Feynman, birth-year, 1918 Heisenberg, birth-year, 1901 to construct an input “Feynman 1918 Heisenberg 1901 Einstein”. By providing such zero-prompt in-context examples to the model, we expect to communicate the underlying relation between subjects and objects. To correctly extrapolate the pattern, the model needs to retrieve Einstein’s birth-year as the completion of the sequence.
More formally, given a training dataset of facts for relation , as well as a test fact , we leverage ICL to construct prompts that elicit information about as
We use to pick facts from and concatenate the tokens corresponding to the subjects and objects, but do not include any other information about . We use space “ ” as the separator token and discuss this choice in detail in Section 4.1. We discuss other design choices for the construction of ZP-LKE in Section 3. When further details are not needed, we simply refer to some input as .
ZP-LKE design satisfies all our design principles (DPs):
DP1: by construction, zero-prompting eliminates prompt hacking and thus, risks of over-fitting and side-channels.
DP2: it only relies on the in-context learning abilities and not meta-linguistic judgments of an LLM.
DP3: by construction, input is LLM-agnostic and hence, enables cross-LLM latent knowledge comparisons.
2.3. Evaluating model outputs
We evaluate the output of model for input in two ways: (1) Open-ended generation that lets the model generate till tokens (Kalo and Fichtel, [n. d.]; Yu et al., 2024) after which the presence of the ground truth is checked within the response, (2) Multiple-choice test that forces the model to predict from a list of options (Jiang et al., 2020).
(1) Response testing in open-ended generation. Given a fact and a model , we provide the input to the model and let it generate for tokens . We consider the answer to be correct if leading to the prediction .
(2) Multiple-choice testing. In the multiple-choice testing, we extract the answer based on the probabilities assigned to the tokens of the corresponding object . To allow for objects consisting of multiple tokens and to be independent of the specific tokenization scheme or LLM (DP3), we compute the object probability over multiple tokens as follows:
where denotes the number of tokens in and is the conditional probability of predicting the -th token of given the preceding tokens , and . To determine whether model knows a fact , we test whether given an input , can choose the correct object from among a set of unique alternatives. Specifically, given fact , we redefine it as , where is a set of plausible but incorrect alternatives. We discuss the choice of in Section 4.
denotes the prediction of for the fact . The predicted object has the maximal object probability within .
We evaluate the factual knowledge of model over a dataset of test facts using accuracy as a metric for both the response test and multiple-choice test:
where is the indicator function.
3. Exploring the design space of ZP-LKE
Our ZP-LKE design avoids many of the reliability concerns of prior works (Jiang et al., 2020; Kalo and Fichtel, [n. d.]; Yu et al., 2024; Hogan et al., 2021; Cao et al., 2021; Petroni et al., 2019). However, ZP-LKE also introduce a few design choices for the input i.e., in Equation (1). One must decide the right , the number of in-context examples included in . Furthermore, it is unclear how ZP-LKE would be affected if some chosen examples are unknown to the model or are incorrect or appeared in a different order.
We study this by varying and introducing unknown or incorrect examples within these examples. While many prior works investigated the number of in-context examples needed for various tasks (Brown et al., 2020; Agarwal et al., 2024; Chen et al., 2023; Lin and Lee, 2024; Pan, 2023), it is worth re-examining them for ZP-LKE for three reasons: (i) prior works report differing results, with some reporting that increasing the number of examples improves performance (Agarwal et al., 2024), while others argue the opposite (Lin and Lee, 2024), (ii) most don’t carefully distinguish between the two learning modes of in-context learning (as noted in Section 1, ZP-LKE relies on one of the modes), and (iii) only a few (Min et al., 2022; Chen et al., 2023) studied the influence of incorrect examples and none studied the impact of unknown examples.
Our experiments help us understand the number of in-context examples needed, as well as how the in-context example’s generation probability changes with different types of noise. We perform an in-depth empirical analysis on a Nobel Laureate dataset for the relation ‘birth year’ (details in Appendix A.1). The dataset consists of facts formatted as .
The number of required in-context samples for communicating both the question and answer format varies across LLMs. In Figure 3, we report multiple choice accuracy (Eq. (4)) for different LLMs evaluated on 900 test samples, with varying numbers of in-context examples () that are randomly sampled from a separate training set using 5 random seeds. As the number of in-context examples increases, the mean accuracy rises while the standard deviation decreases across different LLMs, indicating that the models gradually converge to stable performance.
The dashed vertical lines show the minimum number of examples required by different LLMs to achieve 95% of the accuracy reached with 50 in-context examples. Interestingly, LLMs with higher estimation accuracy require fewer in-context examples than those with lower accuracy to effectively interpret the underlying question. This maybe attributed to the amount of internal knowledge contained in the LLMs. To enable ZP-LKE across all the LLMs, we set for the following experiments.
We delve deeper to further investigate which individual facts may be known or unknown to a model. We examine the generation probability of in-context objects in 200 correct subject ()-object () pairs using the Mistral-7B model. We can see in Figure 4(a)that Mistral-7B model shows a gradual increase in the probability of generating correct objects from left to right on the x-axis (where points on the right have more context to leverage) stabilizing at a mean probability of approximately 0.85. Some objects at later positions, however, have a lower generation probability, suggesting that the LLM may be less confident about its knowledge of the facts corresponding to those pairs. Thus, we can leverage the generation probability as a signal of the LLM’s confidence when evaluating LKEs (see Appendix D). Similar results for additional models are presented in Appendix E.
Models are robust to unknown examples. Next, we investigate the robustness of estimates to the occurrence of unknown examples. We insert unknown examples in two distinct ways: randomly distributed throughout and in a more extreme scenario, where a continuous block of examples is replaced with unknown ones. We selected 40 out of the 200 examples and replaced them with unknown examples created using fictitious names and birth years 222generated via https://en.namefake.com/api. Our findings are shown in Figures 4(b) and 4(c) for distributed and continuous replacement respectively. Unknown examples are marked by red dots, examples immediately following unknown ones in cyan dots and the rest in blue dots. The unknown examples show generation probabilities close to zero, confirming the LLM’s tendency to assign low probabilities to unknown data. However, interestingly, unknown examples minimally impact the generation probability of the surrounding data in both settings.
Models are vulnerable to incorrect examples. Similar to the setup for unknown examples, we also insert 40 (out of 200) incorrect examples randomly (Figure 4(d)) and simultaneously (Figure 4(e)). In our experiments, these incorrect examples are created by altering the birth years of known Nobel laureates and are marked by red dots in the plots. In contrast to inserting unknown examples, the LLM significantly struggles with the injection of incorrect examples. It detrimentally affects the LLM’s performance in both settings thus revealing the vulnerability of the models towards incorrect examples. We highlight one randomly marked yellow star example in Figure 4(a), Figure 4(b), and Figure 4(d) to show how the presence of incorrect samples significantly brings down the probability of the neighboring points.
Summary: The key takeaways while exploring the design space of ZP-LKE are - (a) Different LLMs take varying numbers of in-context samples to comprehend both the question and format of the answer alongside, with being an optimal number for our setup. (b) The models are robust to unknown examples but vulnerable to incorrect examples. To the best of our knowledge, we are the first to distinguish between examples unknown to the model and incorrect examples known to the model and study their impact on in-context learning. As ZP-LKE relies on many examples, this distinction and understanding is important in practice. Also, while (Agarwal et al., 2024) found that the order of examples has varying effects in different domains, we identify the distribution of unknown and incorrect examples as a crucial underlying factor.
4. Experiments and Results
As ZP-LKE inputs are model-agnostic and easy to adapt for a large variety of relations, it can be used to very effectively to conduct cross-LLM latent knowledge comparisons. We leverage ZP-LKE to estimate latent knowledge across 49 open-source (pre-trained and fine-tuned) LLMs, spanning different LLM families (Llama (2), Mistral, Mixtral, Gemma, Falcon, Pythia, Bloom, and OPT) and sizes (from 70M to 822B). To the best of our knowledge, we are the first to evaluate a knowledge estimation framework across a large number of models. We list the models and their simplified names used in this paper in Appendix G, Table 7, and provide a leaderboard of models based on ZP-LKE in Appendix G, Table LABEL:table:model_order. We hope our results and the framework can help future LLM developers reliably and efficiently estimate the latent knowledge of their models.
Dataset: We evaluate the knowledge of models on a large set of facts from the T-REx dataset (Elsahar et al., 2018). We selected relations from T-REx with at least 500 samples that are linked to a minimum of 100 unique objects. We create a list of multiple choices for each sample and ensure that instances with multiple correct objects do not have any of their correct answers in their multiple-choice list. This filtering leads to 50 distinct relations spanning categories like birth dates, directorial roles, parental relationships, and educational lineage. The resulting T-REx Multiple Choice (T-REx-MC) dataset comprises 5,000 training and 20,000 test facts. Appendix A contains detailed information on the dataset and relations.
Choosing the set & its impact on test difficulty: For each fact subject (), relation (), object (), we generate alternative objects to create multiple choices. Note that the alternative objects in are viable choices and cannot be easily eliminated. Therefore, for each fact we select from other facts in the dataset that share the same relationship . For computational feasibility, we sample alternative objects per fact, so that a random guess between has a probability of being correct.
4.1. ZP-LKE vs. prompt-based approaches
We compare the performance of ZP-LKE with the existing prompt-based approaches (Jiang et al., 2020) using both the response accuracy and the multiple-choice accuracy defined in Section 2.2.
ZP-LKE outperforms prompt-based approaches. We randomly sample three human-generated prompts (HGP) and machine-mined prompts (MMP) from (Jiang et al., 2020) for 12 common relations between T-REx-MC and (Jiang et al., 2020). We show that ZP-LKE outperforms HGP and MMP in terms of the accuracy measures by a large margin, across different models and 12 relations in Figure 5; the detailed accuracy for each relation can be found in the Appendix G.2, Figure 14 and 15.
Figure 5(a) shows that ZP-LKE improves the fraction of facts accurately extracted from four open-source models by an average of 35% for HGPs (from 0.45 to 0.61) and 90% for MMPs (from 0.32 to 0.61). In the case of multiple-choice accuracy, having controlled the influence of the answer format, we observe that all the knowledge estimation methods improve in their performance. Alongside, ZP-LKE still outperforms existing approaches by an average of 9.41% for HGPs (from 0.71 to 0.78) and 57% for MMPs (from 0.50 to 0.78). The multiple-choice accuracy metric disentangles the answer format from the question leading to better factual knowledge estimation across the board. Hence, we primarily report the multiple-choice accuracy metric for the experiments in the rest of the paper.
ZP-LKE performs better than FS-LKE with the same number of examples. We adapt ZP-LKE by replacing the separator token “ ” between subjects and objects with three prompts each from HGPs and MMPs for the relation ‘original broadcaster’ and report the multiple choice accuracy in Figure 6. We intend to understand whether the additional prompt tokens can help in communicating the question better. The ZP-LKE with “ ” token performs equally or even better for some models compared to the semantically meaningful prompts from HGP and MMP, which now correspond to the FS-LKE. Thus, relation-specific separators (or prompts) have limited impact on factual knowledge estimation if subject-object pairs are correctly presented. Additionally, finding relation-specific prompts requires hand-crafted efforts or additional computation (Shin et al., 2020), unlike our zero-prompt many-shot approach using (subject, object) pairs. Therefore, ZP-LKE can potentially extend to any fact from knowledge graphs over any LLM, while HGPs and MMPs require additional supervision and relation-specific validation.
4.2. Evaluating Diverse Models and Relations
We investigate the performance of 35 pre-trained LLMs and 14 fine-tuned LLMs across 50 relations using the ZP-LKE framework. Our analysis aims to uncover nuanced insights into the knowledge levels within these models. We will examine the results through two primary lenses: (1) the variations in knowledge across different model families, and (2) the influence of model size and fine-tuning within the same model family on their knowledge attributes.
4.2.1. Comparing different LLM families
Some model families are consistently more knowledgeable than the rest. We sort the model families based on the performance of the model closest to 7B parameters 3337B parameters is a good reference point since all model families except GPT-NEO-X have models within a gap of 1B parameters: Mistral-7B, Gemma-7B, Llama-7B, Falcon-7B, MPT-7B, OPT-6.7B, GPT-J-6B, Pythia-6.9B, and Bloom-7.1B., and the models within each family based on average accuracy across 50 relations. Figure 7 shows that Mistral, Llama2, Gemma, and Llama families have higher performance on most of the relations than Pythia, Bloom, and OPT, indicating the latter’s lower factual knowledge.
Different model families align in their relative factual knowledge. Although different model families have different knowledge levels, they have similar knowledge structures. We investigate the correlations between each model pair’s performance over 50 relations to assess the agreement in their knowledge levels. We compute the average correlations within each model family (e.g. Llama2 7B, 13B, 70B) in Figure 8. Despite differences in architecture and training datasets among model families, there is a significant consensus (correlation ¿ 0.6) regarding the hierarchy of knowledge across various relations. We also compile the three best and worst-performing relations for each model in Table 10, illustrating the consensus among all models. The consistent underperformance across specific relations also suggests that certain types of knowledge are universally less well-represented across different models, regardless of their architecture or size. This consistency in less-known knowledge across models highlights a potential vulnerability that could be exploited if these weaknesses are not addressed. Figure 16 shows the correlations between all the models within each family.
4.2.2. Comparing within the same LLM family
Larger models embed more knowledge with certain exceptions. Figure 7 shows that within each model family, larger models (e.g., Llama-65B) generally outperform smaller ones (e.g., Llama-13B). Models within the same family are typically pre-trained on the same datasets (Biderman et al., 2023; Zhang et al., 2022; Touvron et al., 2023). The results suggest that, when trained on identical datasets, larger models in general capture a broader set of facts. An exception lies in the OPT group of models (Zhang et al., 2022) that have also been trained on the Pile dataset (Gao et al., 2020) like some of the other models (Biderman et al., 2023). This may call for investigating the ways of knowledge injection and validating if that can be attributed to such a performance deviation.
Despite being trained on the same data, models might remember different facts. From the above results, it is not clear if the larger models are subsuming smaller models in their factual knowledge, i.e., do the larger models correctly identify the facts that the smaller models are correct on? To assess this, we compute the subsumption rate :
that measures how much of the fraction of facts from known by smaller model are also recognised by the larger model . A subsumption rate of 1 indicates that all of the smaller model’s knowledge is also contained in the larger model.
| Smallest Model | Largest Model | ||||
| Family | #Parameters | Accuracy | #Parameters | Accuracy | |
| Llama | 7B | 0.699 | 65B | 0.836 | 0.769 |
| Llama-2 | 7B | 0.741 | 70B | 0.846 | 0.801 |
| Gemma | 2B | 0.666 | 7B | 0.750 | 0.710 |
| OPT | 125m | 0.430 | 30B | 0.588 | 0.481 |
| Pythia | 70m | 0.334 | 12B | 0.648 | 0.403 |
| Bloom | 560m | 0.410 | 7.1B | 0.548 | 0.498 |
Table 2 shows the average subsumption rate () between the largest and smallest models in a family, as well as the average accuracy, over all relations for different model families. Interestingly, is relatively low (¡ 0.5) for OPT, Pythia and Bloom (i.e., the larger models know less than 50% of what the smaller models know) and only reaching up to 0.8 for Gemma, Llama and Llama-2. Therefore, even though models within each family are trained on the same datasets and generally agree on the relative knowledge of different relations (Figure 8), there are differences in the knowledge of specific facts they retain from their training data. These discrepancies suggest that simply increasing the model size may not be sufficient to enhance factual knowledge, thus requiring the need for proper factual knowledge injection into the models.
Instruction fine-tuning reduces latent knowledge. Finally, we investigate the effects of chat-based instruction fine-tuning on the factual knowledge of models. Base language models are often fine-tuned (using a mix of supervised and reinforcement learning (Ouyang et al., 2022)) to improve their ability to follow instructions. While previous studies have shown that fine-tuning enhances performance on various benchmarks, its impact on latent knowledge is unclear. Figure 9 illustrates the comparative accuracy of pre-trained models and their fine-tuned counterparts. In almost all cases, the fine-tuned models obtain lower accuracy than their base versions, suggesting that fine-tuning reduces the amount of latent knowledge estimation. A similar observation was made by (Yu et al., 2024). To further assess if fine-tuned models acquire new knowledge, we compute the subsumption rate between pre-trained and fine-tuned versions (Table 11). We find that most latent knowledge in fine-tuned models is already present in base models (high ). This outcome highlights the need for caution when fine-tuning models, as these adjustments might inadvertently compromise with the existing internal knowledge.
5. Concluding Discussion
In this work, we investigate a new way to estimate latent factual knowledge from an LLM. Unlike prior approaches, our method does not engineer prompts (zero-prompting). Rather it relies on LLMs’ in-context learning ability to infer the factual knowledge question and the expected answer format. Our method not only addresses many reliability concerns with prompting, but it also recollects significantly more factual knowledge than prompting. In contrast to prompting, which requires relationship-specific and LLM-specific prompt engineering, Our method can be applied with minimal effort to test factual knowledge of relations across a variety of structured knowledge bases and LLMs. This ability enables us to compare the latent knowledge captured by many different families of open-source LLMs; we expect our results to be of interest to the designers of these LLMs. Finally, to design our zero-prompt many-shot LKE, we explore the impact of the number and order of correct, incorrect, and unknown examples used as inputs; our findings may be of independent interest to developing a better understanding of different learning modes of in-context learning.
A fundamental question posed by our and prior work on estimating latent knowledge in LLMs: What does it mean for an LLM to know a fact? Suppose we tried to infer if an LLM knows the capital of Germany using the input ”France Paris; Spain Madrid; Germany ” and suppose the answer was Berlin. What we have learned is that the LLM knows that the relationship between Germany and Berlin is similar to that between France and Paris or Spain and Madrid. What we have not learned is whether the LLM knows that the relation is called ”capital” in English or ”hauptstadt” in German. The latter is revealed by prompts such as ”The capital of Germany is ”. But, such prompts don’t reveal whether the LLM knows that what Berlin means to Germany is similar to what Paris means to France.
Is one type of knowing facts better than another? It is difficult to answer in general. Neither type of knowing guarantees that the knowledge can be put to use in different contexts and tasks, such as when we ask the LLM where the parliament of Germany is located. However, they lead to different strategies for getting LLMs to generate correct outputs. With the first type of knowing, we can use a list of facts as input such as ”The parliament of France is in Paris; The parliament of Spain is in Madrid; The parliament of Germany is in ”. With the second type of knowing, we can hope to use a chain of thought prompts such as ”The parliament of a country is in its capital. The parliament of Germany is in ”. Nevertheless, one clear takeaway from our study is related to how factual knowledge is latently embedded in an LLM. We show that more factual knowledge can be recollected using in-context learning, i.e., the representations of subjects and objects that share the same relationship, than by prompting with the name of their relationship.
6. Ethical Considerations
Our research utilizes public datasets and open-source LLMs, which mitigates immediate privacy concerns. However, our findings on the factual knowledge capabilities of various LLMs could influence their deployment in real-world applications, potentially leading to over-reliance on models for tasks requiring factual accuracy. We encourage users of our methodology to consider these implications and to use the knowledge estimation techniques responsibly, with appropriate safeguards against potential misuse. Furthermore, as our work may reveal biases or gaps in the factual knowledge of LLMs, we urge developers to address these issues to ensure fair and equitable AI systems.
Appendix A Dataset
A.1. Creation of Nobel laureates dataset from Wikidata
The Nobel Dataset is a collection of biographical information about all Nobel laureates up until the year 2022, totaling 954 individuals. This dataset was curated using data obtained from Wikidata’s querying service444https://query.wikidata.org/. The following attributes are included for each laureate:
Name: The full name of the Nobel laureate.
Birth Year: The year in which the laureate was born.
Award Year: The year(s) in which the laureate was awarded the Nobel Prize.
Nature of Award: A brief description of the reason for the award, including the field of the Nobel Prize (e.g., Physics, Peace).
Gender: The gender of the laureate.
Here are some examples from the Nobel Dataset:
| Name | Birth Year | Award Year | Nature of Award | Gender |
| Albert Einstein | 1879 | 1921 | Physics | male |
| Louis de Broglie | 1892 | 1929 | Physics | male |
| Carl D. Anderson | 1905 | 1936 | Physics | male |
| Polykarp Kusch | 1911 | 1955 | Physics | male |
| Melvin Schwartz | 1932 | 1988 | Physics | male |
| Jerome I. Friedman | 1930 | 1990 | Physics | male |
A.2. Creation of multiple choices from T-REx: TREx-MC
T-REx (Elsahar et al., 2018) is a large-scale alignment dataset that aligns between Wikipedia abstracts and Wikipedia triples. We have utilized the processed version of T-REx available on HuggingFace 555https://huggingface.co/datasets/relbert/t_rex for our experiments. We filtered out the relations that have more than 500 facts and 100 unique object entities. The unique objects ensure having 100 feasible multiple choices for each fact in each relation. We also manually filtered out relations with multiple correct objects (e.g. “America”, “USA”, “American”) to avoid ambiguity. Additionally for relations that have objects in the form of partial matches (e.g. “French”, “French language”), the respective objects have been standardized to uniform values (e.g. “French”). We curated 50 relations for our dataset TREx-MC that essentially consists of subject, relation, multiple choices . The multiple choices comprise the correct answer along with 99 other potential choices. We list the 50 relations in Table 4 below.
The following attributes are included in TREx-MC dataset for each relation:
Subject : The subject entity for each fact.
Object: The object entity or the correct answer for each fact.
Multiple choices: The list of other potential choices for each fact.
Title : The Wikipedia title for each fact.
Text: The Wikipedia abstract corresponding to each fact.
director | father | spouse | child | sibling | composer |
student of | ||||||||||||||||||
| instance of |
genre |
performer | capital | producer | ||||||||||||||||||
| is made by |
developer | publisher |
manufacturer |
| mother |
creates |
inception |
author |
characters | lyrics by | distributed by | home venue |
Some examples from the T-REx-MC dataset for 2 relations are listed in Table 5
| Subject | Object | Multiple choices | Title | Text | ||||||||
| Date of birth | ||||||||||||
| Giovanni Bia | 24 October 1968 |
Giovanni Bia |
| Brian May | 19 July 1947 |
Brian May |
| Composer | ||||||||||||
| Mexico Trilogy | Robert Rodriguez |
Mexico Trilogy |
| Chelsea Walls | Jeff Tweedy |
Chelsea Walls |
Appendix B Inference Setup
We experiment with and use three different inference setups:
Transformers Based Setup: This setup utilizes the utilities present in the transformers library (Wolf et al., 2020) to obtain the log probabilities for generating the different options.
vLLM Based Setup: vLLM ((Kwon et al., 2023)) is a fast inference library for large language models (LLMs). It efficiently manages attention key and value memory using PagedAttention. We observed considerable speed boosts for all 3 LKEs compared to the standard Transformers API.
SGLang Based Setup: SGLang (Zheng et al., 2023) is a structured generation language designed for large language models (LLMs). It speeds up LLM interactions and provides enhanced control through tight integration of its frontend language and backend runtime system. SGLang also leverages Radix Attention to cache common components across queries in the KV cache, enabling substantial speedups. We observed sizable speed boosts for ZP-LKE over vLLM. However, we are constrained by SGLang’s limited model family support at the moment, and only utilize it for the Llama, Mistral, and Mixtral families.
Appendix C Implementation Details
ZP-LKE leverages 50 randomly chosen samples from the training data as in-context examples but does not use the relation name. The base prompt is now composed of 50 different examples followed by the name of the entity being tested. A sample would be “Albert Einstein 14 March 1879 Ernest Rutherford 30 August 1871 … J.J. Thomson 18 December 1856 Max Planck.”
A single forward pass is conducted for each sequence, generating log probabilities for the entire sequence. The common part, represented by the tokens for the base prompt is then removed from the tokens of the concatenated base prompt and option resulting in the log probabilities for the option. If the option is tokenized into multiple tokens, a single probability value is obtained by multiplying the individual token probabilities. The resulting values are normalized across multiple choices, and the option with the highest probability is selected as the correct answer. We use the vLLM Based & SGLang Based Setup for this LKE.
Appendix D Different Metrics
The evaluation metric can readily be adapted to existing classification metrics. For example, we introduced the metric Accuracy@K, a calibrated measure that assesses a model’s confidence in its predictions. This metric quantifies how accurately the model identifies knowledge at specified confidence levels for a given relation. We filter the instances that have their confidence levels ¿ threshold and form the set . Following this, we use our accuracy measure to compute Accuracy@K for varying values of , the results of which are shown in Figure 10.
Appendix E Probabilities of objects in sequence
We first consider correct examples (subject-object pairs) and report the absolute generation probability of objects in corresponding examples. We showed the results for Llama2-7B, Falcon-7B, and Pythia-12B in Figure 11, Figure 12 and Figure 13. Figure 11(a), Figure 12(a) and Figure 13(a) illustrates the probability of each object at various sequence positions; Figure 11(b), Figure 12(b) and Figure 13(b) shows the impact on probabilities after substituting 40 objects dispersed within the sequence with incorrect ones. Figure 11(c), Figure 12(c) and Figure 13(c) visualizes the effect of replacing objects at simultaneous positions. Figure 11(d), Figure 12(d), Figures 13(d), Figure 11(e), Figure 12(e) and Figure 13(e) present the outcomes of using unknown subject-object pairs as replacements. We used a horizontal dashed line showing an average probability of the correct examples. The yellow star marks the example at position 114 in the sequence.
Appendix F Details about the human-generated prompts and machine-mined prompts
We list the used human-generated and machine-mined prompts from (Jiang et al., 2020) in Table 6 with subjects denoted as ¡‘subject’¿.
| Relation Name | Index | HGP Template | MMP Template |
| 1 | {subject} means | {subject} is a small | |
| Instance of | 2 | {subject} is one | {subject} and liberal |
| 3 | {subject} is a | {subject} artist | |
| 1 | {subject} is playing music | {subject} series of | |
| Genre | 2 | {subject} play | {subject} favorite |
| 3 | {subject} performs | {subject} is an american | |
| 1 | {subject} plays in position | {subject} substitutions : | |
| Position played on team / speciality | 2 | {subject} plays at position | {subject} substitutes : |
| 3 | {subject} is in the position | ||
| 1 | The original language of {subject} is | {subject} a. r. rahman | |
| Original language of film/TV show | 2 | The source language of {subject} is | |
| 3 | The default language of {subject} is | ||
| 1 | The capital of {subject} is | {subject} united states embassy in | |
| Capital | 2 | The capital city of {subject} is | {subject} representative legislature |
| 3 | Its capital {subject} is | {subject} rock band from | |
| 1 | {subject} is a native language of | {subject} descent | |
| Native language | 2 | The mother tongue of {subject} is | {subject} speak the |
| 3 | {subject} means | {subject} population or a widely spoken | |
| 1 | {subject} is named after | {subject} and produces | |
| Named after | 2 | {subject} is named for | {subject} variety of standard ) |
| 3 | {subject} is called after | {subject} official | |
| 1 | The official language {subject} is | {subject} professor of | |
| Official language | 2 | {subject} is | {subject} is the official language in |
| 3 | {subject} is officially | {subject} is the official language spoken in | |
| 1 | {subject} is developed by | {subject} was developed by | |
| Developer | 2 | {subject} is created by | {subject} 2008 |
| 2 | {subject} is designed by | {subject} references external links | |
| 1 | {subject} was originally aired on | {subject} premiered on | |
| Original broadcaster | 2 | {subject} was originally broadcast on | {subject} aired on |
| 3 | {subject} was originally shown in | {subject} 2021 | |
| 1 | in order to communicate in | ||
| Language spoken, written or signed | 2 | {subject} used to communicate in | |
| 3 | |||
| 1 | {subject} is represented by music label | {subject} attributed to the | |
| Manufacturer | 2 | {subject} is represented by the record label | {subject} 113 |
| 3 | {subject} is represented by | {subject} cedar point |
Appendix G Additional results
G.1. Model Name Simplification
We list all the models and their simplified names we evaluated in the paper in Table 7.
| Original Name | Simplified Name in Paper |
| mistral-mixtral-8x7B-v0.1 | Mixtral-8x7B |
| Nous-Hermes-2-Mixtral-8x7B-SFT | Mixtral-8x7B-FT1 |
| Nous-Hermes-2-Mixtral-8x7B-DPO | Mixtral-8x7B-FT2 |
| mistral-7b | Mistral-7B |
| mistral-instruct-7b | Mistral-7B-FT1 |
| openhermes-2.5-mistral-7b | Mistral-7B-FT2 |
| llama2-70b | Llama2-70B |
| llama2-70b-chat | Llama2-70B-FT1 |
| llama2-13b | Llama2-13B |
| llama2-13b-chat | Llama2-13B-FT1 |
| vicuna-13b-v1.5 | Llama2-13B-FT2 |
| llama2-7b | Llama2-7B |
| llama2-7b-chat | Llama2-7B-FT1 |
| vicuna-7b-v1.5 | Llama2-7B-FT2 |
| gemma-7b | Gemma-7B |
| gemma-7b-it | Gemma-7B-FT1 |
| gemma-2b | Gemma-2B |
| gemma-2b-it | Gemma-2B-FT1 |
| llama-65b | Llama-65B |
| llama-33b | Llama-33B |
| llama-13b | Llama-13B |
| vicuna-13b-1.3 | Llama-13B-FT1 |
| llama-7b | Llama-7B |
| vicuna-7b-1.3 | Llama-7B-FT1 |
| falcon-7b | Falcon-7B |
| falcon-instruct-7b | Falcon-7B-FT1 |
| mpt-7b | MPT-7B |
| gpt-neox-20b | GPT-NEOX-20B |
| opt-30b | OPT-30B |
| opt-13b | OPT-13B |
| opt-6.7b | OPT-6.7B |
| opt-2.7b | OPT-2.7B |
| opt-1.3b | OPT-1.3B |
| opt-350m | OPT-350M |
| opt-125m | OPT-125M |
| gpt-j-6b | GPT-J-6B |
| pythia-12b | Pythia-12B |
| pythia-6.9b | Pythia-6.9B |
| pythia-2.8b | Pythia-2.8B |
| pythia-1.4b | Pythia-1.4B |
| pythia-1b | Pythia-1B |
| pythia-410m | Pythia-410M |
| pythia-160m | Pythia-160M |
| pythia-70m | Pythia-70M |
| bloom-7.1b | Bloom-7.1B |
| bloom-3b | Bloom-3B |
| bloom-1.7b | Bloom-1.7B |
| bloom-1.1b | Bloom-1.1B |
| bloom-560m | Bloom-560M |
G.2. Additional results on baseline comparison
We compare ZP-LKE on 12 relations from T-REx-MC: capital, named after, developer, manufacturer, genre, instance of, native language, original broadcaster, language spoken written or signed, original language of film / TV show, official language, position played on team/speciality. We chose those 12 relations from T-REx-MC that are found to be in common with (Jiang et al., 2020) where they define the templates for HGP and MMP. We evaluated 4 models (Mistral-7B, Llama-7B, Falcon-7B, and Pythia-12B) and showed all the results in Figure 14 and Figure 15.
G.3. Full order of models and relations
We evaluated 49 models on 50 relations by our ZP-LKE . Table LABEL:table:model_order shows the ordered models by the average accuracy of all the 50 relations. Table 9 shows the ordered relations by the average accuracy of all the 49 models.
| Model | Average Accuracy | Standard Deviation |
| Llama2-70B | 0.8511 | 0.17591 |
| Mixtral-8x7B-SFT | 0.84765 | 0.16919 |
| Mixtral-8x7B | 0.84605 | 0.16653 |
| Llama-65B | 0.84185 | 0.17528 |
| Mixtral-8x7B-DPO | 0.81535 | 0.17580 |
| Llama-33B | 0.81255 | 0.19088 |
| Mistral-7B | 0.79310 | 0.20000 |
| Llama2-13B | 0.78692 | 0.21892 |
| Llama-13B | 0.76845 | 0.21796 |
| Llama2-70B-chat | 0.75815 | 0.21272 |
| Llama2-7B | 0.74945 | 0.24069 |
| Vicuna-13B | 0.74940 | 0.21427 |
| Gemma-7B | 0.74717 | 0.25668 |
| Openhermes-2.5 | 0.74365 | 0.21241 |
| Vicuna-13B-2 | 0.74080 | 0.22807 |
| Vicuna-7B-2 | 0.71695 | 0.24016 |
| Falcon-7B | 0.70190 | 0.27052 |
| Vicuna-7B | 0.70155 | 0.24724 |
| Llama2-13B-chat | 0.69387 | 0.22966 |
| Llama-7B | 0.69260 | 0.27912 |
| Gemma-2B | 0.66600 | 0.28627 |
| GPT-NEOX-20B | 0.66145 | 0.30972 |
| Llama2-7B-chat | 0.66130 | 0.24996 |
| Mistral-instruct-7B | 0.66120 | 0.26173 |
| MPT-7B | 0.64545 | 0.30638 |
| Pythia-12B | 0.63325 | 0.32412 |
| OPT-6.7B | 0.62110 | 0.31313 |
| GPT-J-6B | 0.60965 | 0.32319 |
| OPT-13B | 0.60845 | 0.31017 |
| Pythia-6.9B | 0.59185 | 0.32359 |
| Bloom-7.1B | 0.58270 | 0.31404 |
| OPT-30B | 0.57925 | 0.31813 |
| Pythia-2.8B | 0.57580 | 0.32773 |
| Pythia-1.4B | 0.56330 | 0.33600 |
| Gemma-7B-instruct | 0.55327 | 0.30689 |
| OPT-2.7B | 0.55109 | 0.33260 |
| Bloom-3B | 0.54375 | 0.29199 |
| Pythia-1B | 0.54220 | 0.31560 |
| OPT-1.3B | 0.53610 | 0.33335 |
| Bloom-1.1B | 0.51115 | 0.29346 |
| OPT-350M | 0.50735 | 0.30716 |
| Gemma-2B-instruct | 0.49474 | 0.29628 |
| Pythia-410M | 0.47995 | 0.29598 |
| Bloom-1.7B | 0.47660 | 0.29658 |
| OPT-125M | 0.45195 | 0.29330 |
| Bloom-560M | 0.38465 | 0.28747 |
| Pythia-160M | 0.37145 | 0.28505 |
| Pythia-70M | 0.31260 | 0.27404 |
| Falcon-instruct-7B | 0.00605 | 0.01459 |
| Order | Relation | Average Accuracy |
| 1 | publication date | 0.992071428571429 |
| 2 | inception | 0.983214285714286 |
| 3 | point in time | 0.975714285714286 |
| 4 | drafted by | 0.922214285714286 |
| 5 | native language | 0.8825 |
| 6 | production company | 0.873428571428571 |
| 7 | languages spoken, written or signed | 0.865071428571429 |
| 8 | performer | 0.831142857142857 |
| 9 | has played at | 0.826642857142857 |
| 10 | capital | 0.815857142857143 |
| 11 | is made by | 0.815357142857143 |
| 12 | producer | 0.794714285714286 |
| 13 | record label | 0.794571428571429 |
| 14 | named after | 0.791071428571429 |
| 15 | developer | 0.786928571428571 |
| 16 | publisher | 0.7835 |
| 17 | original broadcaster | 0.781214285714286 |
| 18 | cast member | 0.777 |
| 19 | home venue | 0.771714285714286 |
| 20 | has subsidiary | 0.754142857142857 |
| 21 | manufacturer | 0.749928571428571 |
| 22 | screenwriter | 0.732285714285714 |
| 23 | contains the administrative territorial entity | 0.7255 |
| 24 | creates | 0.721214285714286 |
| 25 | official language | 0.709857142857143 |
| 26 | mother | 0.697857142857143 |
| 27 | part of the series | 0.692214285714286 |
| 28 | founded by | 0.684714285714286 |
| 29 | original language of film or TV show | 0.6825 |
| 30 | date of birth | 0.668857142857143 |
| 31 | date of death | 0.641594184576485 |
| 32 | instance of | 0.588990518331226 |
| 33 | position played on team / speciality | 0.537642857142857 |
| 34 | genre | 0.536 |
| 35 | distributed by | 0.522785714285714 |
| 36 | parent taxon | 0.488428571428571 |
| 37 | director | 0.432928571428571 |
| 38 | author | 0.331285714285714 |
| 39 | father | 0.309214285714286 |
| 40 | educated at | 0.306285714285714 |
| 41 | characters | 0.282857142857143 |
| 42 | composer | 0.276785714285714 |
| 43 | child | 0.259142857142857 |
| 44 | lyrics by | 0.258428571428571 |
| 45 | sibling | 0.250285714285714 |
| 46 | spouse | 0.238785714285714 |
| 47 | is a tributary of | 0.212142857142857 |
| 48 | cause of death | 0.206 |
| 49 | discoverer or inventor | 0.173142857142857 |
| 50 | student of | 0.123357142857143 |
G.4. Relation accuracy correlation of all the pre-trained models
In Table 16, we show the Pearson correlation coefficients between each model pair’s performance across the 50 relations.
| Order/Model | Mistral-8x7B | Mistral-7B | Llama2-70B | Llama2-13B | Llama2-7B | Gemma-7B | Gemma-2B |
| 1 | publication date | point in time | point in time | publication date | publication date | point in time | point in time |
| 2 | point in time | date of death | inception | point in time | inception | inception | inception |
| 3 | inception | publication date | publication date | inception | point in time | publication date | publication date |
| 48 | discoverer or inventor | discoverer or inventor | student of | discoverer or inventor | educated at | date of death | position played on team / speciality |
| 49 | cause of death | cause of death | cause of death | cause of death | cause of death | instance of | discoverer or inventor |
| 50 | student of | student of | is a tributary of | student of | student of | date of birth | student of |
| Order/Model | Llama-65B | Llama-33B | Llama-13B | Llama-7B | Falcon-7B | MPT-7B | GPT-NEOX-20B |
| 1 | publication date | publication date | publication date | publication date | point in time | publication date | publication date |
| 2 | point in time | point in time | point in time | point in time | inception | inception | inception |
| 3 | inception | inception | inception | inception | publication date | point in time | date of death |
| 48 | discoverer or inventor | discoverer or inventor | discoverer or inventor | student of | discoverer or inventor | student of | discoverer or inventor |
| 49 | cause of death | cause of death | cause of death | instance of | is a tributary of | is a tributary of | lyrics by |
| 50 | student of | student of | student of | date of birth | student of | educated at | student of |
| Order/Model | OPT-30B | OPT-13B | OPT-6.7B | OPT-2.7B | OPT-1.3B | OPT-350M | OPT-125M |
| 1 | inception | publication date | publication date | inception | publication date | inception | inception |
| 2 | publication date | inception | inception | publication date | inception | publication date | publication date |
| 3 | point in time | point in time | date of death | point in time | drafted by | point in time | point in time |
| 48 | position played on team / speciality | composer | lyrics by | discoverer or inventor | director | is a tributary of | student of |
| 49 | discoverer or inventor | student of | student of | student of | student of | spouse | is a tributary of |
| 50 | director | date of birth | discoverer or inventor | instance of | date of birth | student of | educated at |
| Order/Model | GPT-J-6B | Pythia-12B | Pythia-6.9B | Pythia-2.8B | Pythia-1.4B | Pythia-1B | Pythia-410M |
| 1 | inception | point in time | publication date | publication date | publication date | publication date | publication date |
| 2 | publication date | publication date | inception | inception | inception | inception | inception |
| 3 | point in time | inception | point in time | point in time | date of death | point in time | drafted by |
| 48 | lyrics by | lyrics by | lyrics by | student of | discoverer or inventor | lyrics by | student of |
| 49 | student of | genre | director | date of birth | lyrics by | student of | discoverer or inventor |
| 50 | date of death | director | date of death | lyrics by | director | date of death | is a tributary of |
| Order/Model | Pythia-160M | Pythia-70M | Bloom-7.1B | Bloom-3B | Bloom-1.7B | Bloom-1.1B | Bloom-560M |
| 1 | publication date | publication date | publication date | publication date | publication date | publication date | publication date |
| 2 | point in time | point in time | inception | inception | inception | inception | inception |
| 3 | date of death | native language | date of death | point in time | point in time | date of death | point in time |
| 48 | student of | official language | lyrics by | is a tributary of | screenwriter | is a tributary of | is a tributary of |
| 49 | capital | instance of | student of | spouse | student of | spouse | director |
| 50 | director | date of death | spouse | student of | spouse | student of | student of |
G.5. Impact of finetuning
We also show the results for the average subsumption rate () in Table 11 for base models and fine-tuned models over the relations in T-REx-MC.
| Family | Model Type | Accuracy | Model Type | Accuracy | |
| Llama-7B | Base | 0.699 | FT-1 | 0.693 | 0.779 |
| Llama-13B | Base | 0.770 | FT-1 | 0.735 | 0.854 |
| Llama2-7B | Base | 0.741 | FT-1 | 0.712 | 0.808 |
| Llama2-7B | Base | 0.741 | FT-2 | 0.664 | 0.790 |
| Llama2-13B | Base | 0.771 | FT-1 | 0.748 | 0.831 |
| Llama2-13B | Base | 0.771 | FT-2 | 0.692 | 0.801 |
| Llama2-70B | Base | 0.846 | FT-1 | 0.739 | 0.811 |
| Mistral-7B | Base | 0.793 | FT-1 | 0.639 | 0.793 |
| Mistral-7B | Base | 0.793 | FT-2 | 0.750 | 0.869 |
| Mixtral-7Bx8 | Base | 0.832 | FT-1 | 0.835 | 0.928 |
| Mixtral-7Bx8 | Base | 0.832 | FT-2 | 0.817 | 0.911 |
| Gemma-2B | Base | 0.666 | FT-1 | 0.488 | 0.577 |
| Gemma-7B | Base | 0.749 | FT-1 | 0.511 | 0.557 |